# Rethinking Latent Representations in Behavior Cloning: An Information Bottleneck Approach for Robot Manipulation

### v2

@@ Abstract
Behavior Cloning (BC) is a widely adopted visual imitation learning method in robot manipulation. Current BC approaches often enhance generalization by leveraging large datasets and incorporating additional visual and textual modalities to capture more diverse information. However, these methods overlook whether the learned representations contain redundant information and lack a solid theoretical foundation to guide the learning process. To address these limitations, we adopt an information-theoretic perspective and introduce mutual information to quantify and mitigate redundancy in latent representations. Building on this, we incorporate the Information Bottleneck (IB) principle into BC, which extends the idea of reducing redundancy by providing a structured framework for compressing irrelevant information while preserving task-relevant features. This work presents the first comprehensive study on redundancy in latent representations across various methods, backbones, and experimental settings, while extending the generalizability of the IB to BC. Extensive experiments and analyses on the CortexBench and LIBERO benchmarks demonstrate significant performance improvements with IB, underscoring the importance of reducing input data redundancy and highlighting its practical value for more practical applications. Project Page: BC-IB Website.
@@ 1 Introduction
Why do we need to explore this? Firstly, the inherent challenges of input data redundancy remain largely unexplored in BC for robot manipulation, despite their potential to significantly impact performance. Secondly, most existing methods lack a solid theoretical foundation to guide the learning process. Then, a key question arises: how can we quantify the redundancy in inputs or representations, and how can we effectively reduce it?
@@ 2.1 Behavior Cloning in Robot Manipulation
Behavioral Cloning (BC), first introduced by (Pomerleau, 1991), is a well-known Imitation Learning (IL) algorithm that learns a policy by directly minimizing the discrepancy between the agent’s actions and those of the expert in the demonstration data. To learn more generalizable representations, one class of visual representation learning methods pre-train on large video datasets of robotics or humans, enabling rapid application of the pre-trained encoder to downstream robotic tasks. Notable examples include VC-1 (Majumdar et al., 2023), R3M (Nair et al., 2023), and Voltron (Karamcheti et al., 2023) . Meanwhile, another line of research focuses on training on even more extensive and diverse datasets with larger models, such as Internet-scale visual question answering and robot trajectory data (Brohan et al., 2023), as well as a vast collection of Internet videos (Cheang et al., 2024). Additionally, some methods further enhance generalization by incorporating additional sources of information. These include inferring textual descriptions based on the robot’s current state (Zawalski et al., 2024), leveraging visual trajectories (Wen et al., 2024) and generated images (Tian et al., 2025), and integrating 3D visual information (Goyal et al., 2023). However, these methods have not deeply analyzed the redundancy in learned latent representations, and most also lack a solid theoretical foundation. Thus we extend the Information Bottleneck (IB) principle to BC, addressing this fundamental gap.
@@ 2.2 Information Bottleneck in Robotics
The Information Bottleneck (IB) principle was first proposed in (Tishby et al., 1999) within the context of information theory. Since then, it has been widely applied in deep learning and various downstream tasks to balance the trade-off between representation accuracy and complexity, including classification (Federici et al., 2019), segmentation (Bardera et al., 2009; Lee et al., 2021), and generative tasks (Jeon et al., 2021). In robotics learning, IB has found notable applications in reinforcement learning, where some works maximize the mutual information between the representation and the dynamics or value function, while restricting the information to encourage the encoder to extract only task-relevant features (Kim et al., 2019; Bai et al., 2021; He et al., 2024). In imitation learning, it has been introduced to solve copycat problems (Wen et al., 2020). Different from prior works, we introduce IB into Behavior Cloning to explore and experimentally validate the redundancy in latent representations in robotics. Additionally, we demonstrate its effectiveness through detailed theoretical analyses.
Figure 2: Model architectures used in this study. Based on feature fusion methods, we categorize the BC methods in robot manipulation into two types: spatial fusion and temporal fusion. After extracting features from each modality a), spatial fusion b) extracts spatial features at a given time step or concatenates features across multiple time steps using encoders like MLPs or CNNs. Temporal Fusion c) fuses input features by modeling dynamic relationships and dependencies between time steps using RNNs or Temporal Transformers. The latent representations are then decoded into actions via the policy head.
@@ 4.1 Model Architecture
Before introducing IB, we first need to define the input $X$ and latent representations $Z$. According to current methods in BC for robot manipulation, as discussed in Section 3.1, the input is typically multimodal, meaning it not only includes RGB images but may also incorporate the robot’s proprioceptive state, language instructions, and more. Previous work has shown that proprioceptive states can lead to overfitting (Wang et al., 2024). Additionally, for the sake of convenience in visualizing $I(X,Z)$, we do not treat the image alone as $X$, as done in previous studies. Instead, we use features extracted from all modalities through respective feature extractors as our input $X$, i.e.,
@@
where $\text{Enc}_{(\cdot)}$ denotes the feature extractor of each modality. Then, regarding how to process the input $X$, or how to fuse information from multiple modalities into latent representations $Z$, we categorize the BC methods in robot manipulation based on feature fusion methods into two types: spatial fusion and temporal fusion.
As illustrated in Figure 2 b), spatial fusion involves extracting spatial features from data at a given time step or concatenating features across multiple time steps along the feature dimensions. This approach does not explicitly differentiate between time steps but instead processes the aggregated features as a whole, emphasizing the modeling of inter-feature relationships. The spatial fusion module can be implemented using Multi-Layer Perceptrons (MLPs), Convolutional Neural Networks (CNNs), Spatial Transformers, or even simple concatenation operations. On the other hand, as illustrated in Figure 2 c), temporal fusion integrates input features by capturing dynamic relationships and dependencies across time steps, enabling the modeling of both long-term and short-term temporal dynamics in sequential data. Temporal fusion modules can be implemented using Recurrent Neural Networks (RNNs), Long Short-Term Memory networks (LSTMs), or Temporal Transformers.
The latent representation $Z$, which integrates both spatial and temporal information, is then passed through a policy head to generate actions. Existing policy heads primarily focus on using MLP, Gaussian Mixture Model (GMM), and diffusion-based policy (DP) heads (Chi et al., 2023; Reuss et al., 2024). For simplicity and clearer empirical demonstration, we use an MLP as the policy head.
@@ 4.3 Theoretical Analysis
We provide a theoretical analysis of our BC-IB objective in Equation 5. Theorem 4.1 and Theorem 4.2 reveal that the generalization error admits an upper bound governed by the mutual information between the input and the latent representation. By minimizing this mutual information, we effectively tighten the upper bound, thereby improving the model’s generalization performance. Theorem 4.3 elucidates the optimization challenges associated with complex input states. When the input $O$ contains a large amount of state information, as depicted in Figure 2, directly minimizing the mutual information between $O$ and the latent representation $Z$ becomes computationally impractical. To address this, the input $O$ is first compressed into an intermediate embedding $X$ via a fusion network $f$, and the mutual information between $X$ and $Z$ is minimized instead. The theorem establishes that, under certain conditions, this intermediate approach can approximate the ideal optimization results, provided that the embedding $X$ sufficiently preserves the essential information from the original input $O$.
@@
Let $J^{\varepsilon}=\mathbb{E}_{P_{f^{\varepsilon},\phi^{\varepsilon}}(o,x,z)}[J(z;\theta^{\varepsilon})],\quad J^{\star}=\mathbb{E}_{P_{\phi_{o}^{\star}}(o,z)}[J(z;\theta^{\star})]$.
Assume the mutual information gap satisfies the following, for any $\delta$ have
@@ 5.1 Embodied Evaluation
Benchmarks. We mainly evaluate BC with IB across two benchmarks, CortexBench (Majumdar et al., 2023) and LIBERO (Liu et al., 2024). CortexBench is a single-task benchmark. For validation, we selected four imitation learning-related simulators, encompassing a total of 14 tasks: Adroit (2 tasks) (Rajeswaran et al., 2018), Meta-World (5 tasks) (Yu et al., 2020), DMControl (5 tasks) (Tassa et al., 2018), and TriFinger (2 tasks) (Wuthrich et al., 2021). During evaluation, the number of validation trajectories is set to 25, 10, 25, and 25, respectively. LIBERO is a language-conditioned multi-task benchmark. For evaluation, we select four suites: LIBERO-Goal (10 tasks), LIBERO-Object (10 tasks), LIBERO-Spatial (10 tasks), and LIBERO-Long (10 tasks), each focusing on the controlled transfer of knowledge related to task goals, objects, spatial information, and long-horizon tasks, respectively. During evaluation, the number of validation trajectories is set to 20.
@@
Baselines. In CortexBench, we evaluate four representation learning models: R3M (Nair et al., 2023), Voltron (Karamcheti et al., 2023), VC-1 (Majumdar et al., 2023), and MPI (Zeng et al., 2024). In line with the original papers, we use the pre-trained versions of these models to facilitate their application to downstream tasks, keeping the image encoders frozen. Additionally, we introduce two full fine-tuning baselines by replacing these encoders with part of uninitialized ResNet-18 (He et al., 2016) and ViT-S (Dosovitskiy, 2021), denoted as ResNet and ViT, respectively. All methods employ the two fusion techniques described in Section 4.1. For spatial fusion, we use an MLP, and for temporal fusion, we utilize a Temporal Transformer. In LIBERO, we implement four vision-language policy networks. One of them uses a spatial fusion approach, which employs ResNet as the image encoder and an MLP as the fusion module, referred to as BC-MLP. The other three use temporal fusion. Following the original paper, we rename them based on the combination of the image encoder and fusion module: BC-RNN, BC-Transformer, and BC-VILT (Liu et al., 2024). The policy head for all methods is fixed as an MLP. Notably, all baselines with IB are referred to as BC+IB.
@@
Implementation. In CortexBench, for four partial fine-tuning methods, we train for 100 epochs on each task using the Adam optimizer with a learning rate of 1e-3, a batch size of 512, and weight decay of 1e-4, with learning rate decay applied using a cosine annealing schedule. For the two full fine-tuning methods, we train for 50 epochs with a learning rate of 1e-4 and a batch size of 256. All other parameters remain unchanged. In LIBERO, we train for 50 epochs using the AdamW optimizer with a learning rate of 1e-4 and a batch size of 64, decayed using a cosine annealing schedule. For BC+IB methods, the model used in MINE consists of a two-layer MLP, with a learning rate of 1e-5. The Lagrange multiplier in Equation 6 ranges from 1e-4 to 5e-3 in this work.
@@
Model Selection. For single-task benchmark CortexBench, we test the model every 5 or 10 epochs and select the model with the highest success rate. For multi-task benchmark LIBERO, we select the model from the final epoch.
@@
The appendix will provide detailed descriptions of each benchmark (Section B.1), all baselines (Section B.2), implementation details (Section B.3), and the rationale behind the model selection (Section B.4).
@@ 5.2 Performance on Cortexbench
Finding 1: For simple single-task scenarios, spatial fusion is more efficient and effective than temporal fusion. As shown in Figure 3 b), the performance of methods with temporal fusion drops significantly. From Figure 3 a), this can be attributed to the slower loss reduction in methods using temporal fusion, which results in higher loss at the same training epoch. Therefore, we focus exclusively on presenting the results for methods employing spatial fusion.
Finding 4: For simple single-task downstream tasks, full fine-tuning of a simple, uninitialized model (ResNet) is sufficient and may even outperform a pre-trained larger model. However, the latter is more efficient for faster fine-tuning and deployment, and proves to be more effective for more complex tasks (Burns et al., 2023).
Figure 3: a) BC loss variation for ResNet in spatial and temporal fusion methods on the bin-picking task of the Meta-World. b) Averaged success rates of ResNet and VC1 in spatial and temporal fusion methods across the Meta-World and DMControl.

### v3

# Rethinking Latent Redundancy in Behavior Cloning: An Information Bottleneck Approach for Robot Manipulation

@@ Abstract
Behavior Cloning (BC) is a widely adopted visual imitation learning method in robot manipulation. Current BC approaches often enhance generalization by leveraging large datasets and incorporating additional visual and textual modalities to capture more diverse information. However, these methods overlook whether the learned representations contain redundant information and lack a solid theoretical foundation to guide the learning process. To address these limitations, we adopt an information-theoretic perspective and introduce mutual information to quantify and mitigate redundancy in latent representations. Building on this, we incorporate the Information Bottleneck (IB) principle into BC, which extends the idea of reducing redundancy by providing a structured framework for compressing irrelevant information while preserving task-relevant features. This work presents the first comprehensive study on redundancy in latent representations across various methods, backbones, and experimental settings, while extending the generalizability of the IB to BC. Extensive experiments and analyses on the CortexBench and LIBERO benchmarks show consistent performance improvements with IB across various settings, underscoring the importance of reducing input data redundancy and highlighting its practical value for real-world applications. Project Page: BC-IB Website.
@@ 1 Introduction
Why do we need to explore this? Firstly, the inherent challenges of input data redundancy remain largely unexplored in BC for robot manipulation, despite their potential impact on policy performance and generalization. Secondly, most existing methods lack a solid theoretical foundation to guide the learning process. This raises a key question: how can we formally characterize and reduce redundancy in inputs or representations in a theoretically grounded way?
@@ 2.1 Behavior Cloning in Robot Manipulation
Behavioral Cloning (BC), first introduced by (Pomerleau, 1991), is a well-known Imitation Learning (IL) algorithm that learns a policy by directly minimizing the discrepancy between the agent’s actions and those of the expert in the demonstration data. To learn more generalizable representations, one class of visual representation learning methods pre-trains on large video datasets of robotics or humans, enabling rapid application of the pre-trained encoder to downstream robotic tasks. Notable examples include VC-1 (Majumdar et al., 2023), R3M (Nair et al., 2023), and Voltron (Karamcheti et al., 2023) . Meanwhile, another line of research focuses on training on even more extensive and diverse datasets with larger models, such as Internet-scale visual question answering and robot trajectory data (Brohan et al., 2023), as well as a vast collection of Internet videos (Cheang et al., 2024). Additionally, some methods further enhance generalization by incorporating additional sources of information. These include inferring textual descriptions based on the robot’s current state (Zawalski et al., 2024), leveraging visual trajectories (Wen et al., 2024) and generated images (Tian et al., 2025), and integrating 3D visual information (Goyal et al., 2023). However, these methods have not deeply analyzed the redundancy in learned latent representations, and most also lack a solid theoretical foundation. Thus we extend the Information Bottleneck (IB) principle to BC, addressing this fundamental gap.
@@ 2.2 Information Bottleneck in Robotics
The Information Bottleneck (IB) principle was first proposed in (Tishby et al., 1999) within the context of information theory. Since then, it has been widely applied in deep learning and various downstream tasks to balance the trade-off between representation accuracy and complexity, including classification (Federici et al., 2019), segmentation (Bardera et al., 2009; Lee et al., 2021), and generative tasks (Jeon et al., 2021). In robotics learning, IB has found notable applications in reinforcement learning, where some works maximize the mutual information between the representation and the dynamics or value function, while restricting the information to encourage the encoder to extract only task-relevant features (Kim et al., 2019; Bai et al., 2021; He et al., 2024). In imitation learning, it has been introduced to solve copycat problems (Wen et al., 2020). Different from prior works, we introduce IB into Behavior Cloning to explore and empirically validate the redundancy in latent representations in robotics. Additionally, we demonstrate its effectiveness through detailed theoretical analyses.
Figure 2: Model architectures used in this study. Based on the feature fusion strategy, we categorize BC methods for robot manipulation into two types: spatial fusion and temporal fusion. After extracting features from each modality (a), spatial fusion (b) processes spatial features at a single time step or concatenates features across multiple time steps using encoders such as MLPs or CNNs. Temporal fusion (c) captures dynamic relationships and dependencies across time using models like RNNs or Temporal Transformers. The resulting latent representations are then decoded into actions through the policy head.
@@ 4.1 Model Architecture
Before introducing IB, we first define its input and latent representations. Traditional IB methods (Amjad & Geiger, 2019; Pacelli & Majumdar, 2020; Wan et al., 2021) typically apply the bottleneck to a single modality (e.g., images or states) and their corresponding latent features, following the information flow $O\rightarrow Z\rightarrow A$. In contrast, BC for robot manipulation is more complex than earlier control or single-modal tasks, as it requires models to process diverse, multimodal data. This data not only includes RGB images but may also incorporate the robot’s proprioceptive state, language instructions, and other modalities, making effective feature fusion essential. Strictly adhering to the conventional IB paradigm would involve constraining each input modality and its corresponding features separately, resulting in a pipeline that is inelegant, difficult to scale, overly complex, and unable to capture cross-modal associations. Furthermore, previous work has shown that proprioceptive states can lead to overfitting (Wang et al., 2024).
As a result, we do not treat image or other modalities separately as inputs to IB, as done in previous studies. Instead, we concatenate features extracted from all modalities through respective feature extractors as our input $X$, i.e.,
@@
where $\text{Enc}_{(\cdot)}$ denotes the feature extractor of each modality. This results in the information flow $O\rightarrow X\rightarrow Z\rightarrow A$, and brings several advantages: (1) It enables unified redundancy reduction across all modalities. (2) In practice, encoders are often frozen, and this approach is more effective under such conditions. (3) It scales better to a broader range of robotic algorithms. Then, regarding how to process the input $X$, or how to fuse information from multiple modalities into latent representations $Z$, we categorize BC methods in robot manipulation into two types based on their feature fusion strategies: spatial fusion and temporal fusion.
As illustrated in Figure 2 (b), spatial fusion involves extracting spatial features from data at a given time step or concatenating features across multiple time steps along the feature dimensions. This approach does not explicitly differentiate between time steps but instead processes the aggregated features as a whole, emphasizing the modeling of inter-feature relationships. The spatial fusion module can be implemented using Multi-Layer Perceptrons (MLPs), Convolutional Neural Networks (CNNs), Spatial Transformers, or even simple concatenation operations. These methods are primarily designed to learn highly generalizable visual encoders by leveraging large-scale human video datasets. The pretrained encoders are then fine-tuned for downstream robotic tasks (Nair et al., 2023; Karamcheti et al., 2023; Majumdar et al., 2023; Zeng et al., 2024).
On the other hand, as illustrated in Figure 2 (c), temporal fusion integrates input features by capturing dynamic relationships and dependencies across time steps. This enables the modeling of both long-term and short-term temporal dynamics in sequential data. Temporal fusion modules can be implemented using Recurrent Neural Networks (RNNs), Long Short-Term Memory networks (LSTMs), or Temporal Transformers. These methods are commonly incorporated into approaches that utilize Transformer-based backbones (Wu et al., 2024; Li et al., 2024b; Liu et al., 2024).
The latent representation $Z$, which integrates both spatial and temporal information, is then passed through a policy head to generate actions. Existing policy heads primarily focus on using MLPs, Gaussian Mixture Model (GMM), and diffusion-based policy (DP) heads (Chi et al., 2023; Reuss et al., 2024). For simplicity and clearer empirical demonstration, we use an MLP as the policy head.
@@ 4.3 Theoretical Analysis
We provide a theoretical analysis of our BC-IB objective in Equation 5. We adapt Theorem 4.1 and Theorem 4.2 to reveal that the generalization error is upper-bounded by the mutual information between the input $O$ and the latent representation $Z$, following the information flow $O\rightarrow Z\rightarrow A$. Minimizing this mutual information tightens the bound and improves generalization. However, when $O$ is diverse and multimodal, directly minimizing the mutual information between each modality and its corresponding $Z$ is computationally intractable and unnecessarily complex. To address this, we extract and concatenate features from all modalities into an intermediate feature $X$, obtain $Z$ via a fusion network $f$, and minimize the mutual information between $X$ and $Z$ instead. To validate the compatibility of this paradigm with the original theorems, we present Theorem 4.3. The theorem establishes that, even if we optimize $I(X;Z)$ by applying the bottleneck at an intermediate feature level $X$, as long as $X$ preserves the essential structure of the original input $O$, we are effectively controlling $I(O;Z)$, with the difference bounded by a small constant $\delta$.
@@
Let $J^{\varepsilon}=\mathbb{E}_{P_{f^{\varepsilon},\phi^{\varepsilon}}(o,x,z)}[J(z;\theta^{\varepsilon})],\quad J^{\star}=\mathbb{E}_{P_{\phi_{o}^{\star}}(o,z)}[J(z;\theta^{\star})]$. Assume the mutual information gap satisfies the following condition: for any $\delta$, we have
@@ 5.1 Embodied Evaluation
Simulation Benchmarks. We mainly evaluate BC with IB across two benchmarks, CortexBench (Majumdar et al., 2023) and LIBERO (Liu et al., 2024). CortexBench is a single-task benchmark. For validation, we selected four imitation learning-related simulators, encompassing a total of 14 tasks: Adroit (2 tasks) (Rajeswaran et al., 2018), Meta-World (5 tasks) (Yu et al., 2020), DMControl (5 tasks) (Tassa et al., 2018), and TriFinger (2 tasks) (Wuthrich et al., 2021). During evaluation, the number of validation trajectories is set to 25, 10, 25, and 25, respectively. LIBERO is a language-conditioned multi-task benchmark. For evaluation, we select four suites: LIBERO-Goal (10 tasks), LIBERO-Object (10 tasks), LIBERO-Spatial (10 tasks), and LIBERO-Long (10 tasks), each focusing on the controlled transfer of knowledge related to task goals, objects, spatial information, and long-horizon tasks, respectively. During evaluation, the number of validation trajectories is set to 20.
Real-world Evaluation. As shown in Figure 4, our real-world experiments use a 6-DOF UR5 arm equipped with a Robotiq 2F-85 gripper and a RealSense L515 base camera for RGB image capture. Following the simulation setup, we evaluate both a single-task setting and a more challenging language-conditioned multi-task setting. The latter introduces increased distractor objects, randomized object positions, and unseen instances during evaluation to assess generalization. We design two tabletop manipulation tasks: Pick, where the robot lifts an object from the table, and Put (Pick and Place), where the robot picks up an object and places it into a bowl. Demonstrations are collected using a 3D mouse with only the base camera. In the single-task setting, we use 25 demonstrations for Pick and 50 for Pick-and-Place. In the multi-task setting, we collect 800 demonstrations in total, with 200 per task. During evaluation, each task is tested over 10 trajectories.
@@
Baselines. In CortexBench, we evaluate four visual imitation learning models: R3M (Nair et al., 2023), Voltron (Karamcheti et al., 2023), VC-1 (Majumdar et al., 2023), and MPI (Zeng et al., 2024). Following the original papers, we use pre-trained models with frozen image encoders for downstream tasks. Additionally, we introduce two full fine-tuning baselines by replacing the encoders with partially uninitialized ResNet-18 (He et al., 2016) and ViT-S (Dosovitskiy, 2021), denoted as ResNet and ViT, respectively. All methods use the two fusion techniques from Section 4.1: an MLP for spatial fusion and a Temporal Transformer for temporal fusion. In LIBERO, we implement four vision-language policy networks. One of them uses a spatial fusion approach, which employs ResNet as the image encoder and an MLP as the fusion module, referred to as BC-MLP. The other three use temporal fusion. Following the original paper, we rename them based on the combination of the image encoder and fusion module: BC-RNN, BC-Transformer, and BC-VILT (Liu et al., 2024). The policy head for all methods is fixed as an MLP. In real-world evaluation, we adopt VC-1 (Majumdar et al., 2023) for the single-task setting and CogAct (Li et al., 2024a) for the language-conditioned multi-task setting. Notably, all baselines with IB are referred to as BC+IB.
@@
Implementation. In CortexBench, for four partial fine-tuning methods, we train for 100 epochs on each task using the Adam optimizer with a learning rate of 1e-3, a batch size of 512, and weight decay of 1e-4, with learning rate decay applied using a cosine annealing schedule. For two full fine-tuning methods, we train for 50 epochs with a learning rate of 1e-4 and a batch size of 256. In LIBERO, we train for 50 epochs using the AdamW optimizer with a learning rate of 1e-4 and a batch size of 64, decayed using a cosine annealing schedule. In real-world evaluation, we train VC-1 for 200 epochs using the Adam optimizer with a learning rate of 1e-3 and a batch size of 512. CogAct is trained for 8k steps with the AdamW optimizer, using a learning rate of 2e-5 and a batch size of 128. For BC+IB methods, the model used in MINE consists of a two-layer MLP, with a learning rate of 1e-5. The Lagrange multiplier in Equation 6 ranges from 1e-4 to 5e-3 in this work.
@@
Model Selection. For the single-task benchmark CortexBench, we test the model every 5 or 10 epochs and select the model with the highest success rate. For the multi-task benchmark LIBERO, we select the model from the final epoch. For real-world evaluation, we follow the corresponding strategy for each setting as described above.
@@
The appendix provides detailed descriptions of each benchmark (Section B.1), all baselines (Section B.2), implementation details (Section B.3), and the rationale behind the model selection (Section B.4).
Figure 3: (a) BC loss variation for ResNet in spatial and temporal fusion methods on the bin-picking task of the Meta-World. (b) Averaged success rates of ResNet and VC1 in spatial and temporal fusion methods across the Meta-World and DMControl.
@@ 5.2 Performance on Cortexbench
Finding 1: For simple single-task scenarios, spatial fusion is more efficient and effective than temporal fusion. As shown in Figure 3 (b), the performance of methods with temporal fusion drops significantly. From Figure 3 (a), this can be attributed to the slower loss reduction in methods using temporal fusion, which results in higher loss at the same training epoch. Therefore, we focus exclusively on presenting the results for methods employing spatial fusion.
@@
Finding 4: In some benchmarks, particularly Trifinger, the improvement is minimal. We attribute this to the benchmark itself containing very limited redundancy in the visual input, which consists primarily of the robot arm and a single object.
Finding 5: For simple single-task downstream tasks, full fine-tuning of a simple, uninitialized model (ResNet) is sufficient and may even outperform a pre-trained larger model. However, the latter is more efficient for faster fine-tuning and deployment, and proves to be more effective for more complex tasks (Burns et al., 2023).
Figure 4: Real-world robot experiments conducted on a tabletop setup with two settings. (a) Left: the experimental setup. (a) Right: an example of predicted trajectories alongside policy execution. (b) and (c): quantitative evaluation results across two settings, where blue denotes the vanilla BC method and red denotes the method with IB. denotes the unseen tasks.